OpenStudy (anonymous):

f(x) = (x + 2)2 and g(x) = x2 − (x + 2)2. Find f(x) − g(x). A) 2x2 + 8x + 8 B) 3x2 + 8x + 8 C) x2 + 8x + 8 D) 4x + 4 E) x2 − 8x + 8

OpenStudy (nurali):

f(x) − g(x). (x + 2)^2-x^2−(x + 2)^2 x^2+4x+4-x^2+x^2+4x+4 x^2+8x+8

OpenStudy (anonymous):

thank you!

OpenStudy (nurali):

Anytime.

OpenStudy (anonymous):

can you help me on this one f(x)=x^3+6x^2+x^1/2 and g(x)=x^1/2 find f(x)/g(x)

OpenStudy (anonymous):

@Nurali

OpenStudy (nurali):

x^5/2+6x^3/2+1

OpenStudy (anonymous):

thank you so much!

OpenStudy (nurali):

Anytime.